# Characterization of FBK NUV-HD-Cryo SiPMs near LHe temperature

### v3

@@ III.2.1 PDE tests at 10 K
We implemented the photocurrent method to measure the PDE, as introduced in reference [34]. The PDE is calculated as Eq. (1),
where and are the measured current of the SiPMs with the light on and off, respectively; is the gain of the SiPMs; is the charge of an electron; stands for Excess Charge Factor, which represents the fake efficiency prompt contributed from AP and CT [44, 35]; is the incident photons.
@@
where and represent the measured current for the bias voltage above or below the breakdown voltage in the dark, respectively; the definition of and are the same as Eq. (1); is the average number of photoelectrons per second in dark condition, which is equivalent to the average number of the Poisson statistical photoelectrons in a test [40, 41, 33, 42, 43, 34]. As mentioned above in section III.1, the SiPMs’ DCR is as low as 1 Hz near LHe temperature in the dark, which leads to 0.1 pA current. While the measured current is 0.1 nA, essentially the detector system’s dark current. So, we illuminated the SiPMs with a weak pulsed light for the ECF tests to have a A scale current. Effectively, our testing condition of “weak light at LHe temperature” is equivalent to the conventional one of “dark at RT”.
@@
The SiPMs were illuminated with 405 nm or 503 nm light from an LED driven by a DG535 pulse generator. For PDE tests, the pulses from the DG535 are 4 V amplitude, 240 ns width, and 10 kHz frequency. Whenever the DG535 generated a pulse, it also triggered an 8 GHz bandwidth, 25 GSa/s sample rate oscilloscope to record the waveform showing on the screen. A typical single photoelectron (Phe) signal of the FBK SiPMs at 10 K is shown in Fig. 6. Given that (a) the effective time window of the pulses is 2.4e-3 s (= 240 ns 10 kHz) per second, and (b) the DCR is about 1 Hz near LHe temperature, the dark current events only have a chance of 0.24% (= 2.4e-3 s / 1 s) to show up in data. Therefore, the DCR will not smear the PDE results.
@@
As mentioned above, we must figure out the average number of photoelectrons to calculate PDE. We get it through a fit on a charge histogram as shown in Fig. 7. The histogram was obtained by integrating thousands of analog signals in a 10 s window. Fig. 6 shows such a typical signal. During data-taking, the DG535 was tuned to generate weak enough light to produce 1.5 Phe in the SiPMs on average. The oscilloscope was triggered by the DG535. The signals were firstly recorded in the oscilloscope, then analyzed offline with ROOT [45]-based scripts. A recorded raw signal often contained roughly 2000 data points, which were integrated directly (without baseline subtraction), then divided by a 50 ohm (the input resistance of the oscilloscope) to convert into charge. Next, we multiplied the charge with an estimated gain to get the tentative number of Phe, as shown in Eq. (4),
@@
where the is the integral of raw signals, 50 is the input resistance of the oscilloscope, the is the estimated gain of the SiPMs (often is 1.0 106), 1.6 10-19 is the charge of an electron. Fig. 7 shows a typical charge distribution and a fit. The fit function is Gaussian convoluted Poisson, as shown in Eq. (7).
@@
The meaning of the parameters in the above equations are the following. “”, “”, and “” are the fit functions correspond to the pedestal (or 0 Phe), 1 to N Phe peaks, and the whole histogram, respectively. “” is the average number of the Poisson function (“” is also equivalent to the “N_Phe” in Fig. 7 and the in Eq. (3)). “” means the scale factor of the fit function on the Y-axis. “” indicates the shift on the X-axis, which is the analog signals’ baseline shift on the oscilloscope. “” corresponds to the scale factor on the X-axis, which is the “correction” to the estimated gain in Eq. (4). “” and “” are the 1- of the one Phe peak and the pedestal distribution, respectively.
In Fig. 7, the solid green curve represents the fit function, composed of several Poisson-weighed Gaussian functions, and the broken lines marked with “0 Phe”, “1 Phe” …, and “4 Phe” are the mean values of these functions, as the dotted colorful curves show. However, the fit results in the statistics box do not correspond to the green curve; instead, they represent the fit function that has only been applied partly on the histogram ( 2.0 Phe), as the red-on-green part of the curve shows. We limited our fit function within the range to avoid smearing the fit by AP and CT events. The meaning of the parameters in the statistics box are the same as Eq. (5) and (6). The “” is equivalent to the “” in the equations.
As mentioned above, since the gain in Eq. (4) was an estimated value; as a result, the X-axis in Fig. 7 does not represent the exact number of Phe; instead, it’s only a tentative one. With the fit results, the real number of Phe can be obtained with Eq. (8).
With the analysis shown in Eq. (8), the new histogram’s X-axis number would line up with the number of Phe, as we have done in our previous work [46]. However, we did not do the analysis in Fig. 7 because (a) it is not absolutely necessary to do that in some circumstances since as long as the N_Phe (N 1) positions can be located on the histogram’s X-axis precisely, the photosensor’s response would be well understood [47, 48] and (b) the following CT analysis relies on Fig. 7. For more info on the charge histogram and the fit function, please refer to reference [46]. In the specific example, the average number of photoelectrons, 1.405 0.304 Phe, is the ND to be substituted in Eq. (3) for PDE calculation. Fig. 7 does not have separated Phe peaks as the similar SiPMs tested in references [34, 24]. We currently are not able to fully interpret the feature. This topic will be further discussed in section VI.1.1.
@@
All of the parameters we measured or calculated for the PDE test are shown in table 1. Submitting these parameters into Eq. (3) gives the PDE of (39.50 8.77)%. Following the same procedure, the PDE values measured with other bias voltages and 530 nm light can also be obtained, as shown in Fig. 8. From the plot, we can see that for a bias voltage of 10 V OV or more, the PDE is equal to or greater than 40%.
@@
1.410.30 | 5.300.09 | 0.120.07 | 2.760.02 | 0.020.01 | (6.830.31)E6 | 39.50 8.77 |
As shown in Fig. 8, the uncertainties of the PDE under the two wavelengths are around 8%. According to our analysis, 95% of the uncertainty was from the average number of photoelectrons in charge fit, ND. A feature of separated photoelectron peaks, such as the 6 smaller (4.0 4.0 mm2) FBK SiPMs measured in reference [34, 24], should be helpful to decrease ND’s error bar and the PDE’s uncertainty significantly.
III.2.2 PDE tests at RT
Although our SiPMs are supposed to work at LHe temperatures, we measured the PDE at RT with a similar setup and procedure at 10 K, as shown in Fig. 9. The PDE results are consistent with the measurements on the similar FBK SiPMs at RT [34, 24]. From Fig. 8 and 9, we can see (or extrapolate) that the PDE is significantly higher at RT than at 4 K under the same bias voltage. Specifically, for the bias of 5 V OV, the discrepancy between RT and 4 K for 405 nm and 530 nm light are 27% (45% - 18%) and 26% (42% - 16%), respectively. Though a thorough understanding of the feature of the lower the temperature, the lower the SiPMs PDE still lacks, our test results are consistent with references [49, 50].
III.2.3 Gain tests at 10 K
We tested the SiPMs’ gain at 10 K with 405 nm light. The gain calculation was based on the charge histogram as Fig. 7. With the fit results and Eq. (8), a new histogram can be obtained in which the values on the X-axis would line up with the number of Phe (including the single photoelectron peak) and the real gain can be extrapolated, as we have done in our previous work [46]. As discussed above, with the fit results, X_scale, the gain can be “corrected”. Fig. 10 shows the gain under variant OV voltages. With the same bias voltage of OV = 5 V, the gain of the NUV-HD-Cryo SiPMs at 10 K was 0.23 106; while the gain of the NUV-HD-LF type of SiPMs at 40 K was 0.32 106 [27, 24]. Given that the feature of the lower the temperature, the smaller the gain was observed in the two references, we consider our results consistent with the previous measurements.
@@ III.3 DCR, AP, and CT tests at 10 K
The dark current rate (DCR), after-pulse (AP), and cross-talk (CT) are critical parameters for SiPMs. Although these parameters were measured for two types of FBK SiPMs from RT down to 40 K, as shown in reference [34, 27], we must characterize them near LHe temperature again for the following reasons. (a) The SiPMs under our tests are not the same model as the references according to our discussion with FBK engineers [37], though it is similar to the NUV-HD-LF one. (b) The dimension of our SiPMs is 11.7 7.9 mm2, which is a factor of six greater than the ones tested in reference [34, 27], 4.0 4.0 mm2; the cell pitch of our SiPMs is 30 30 m2, which is also bigger than the 25 25 m2 one in the references. (c) For FBK SiPMs, the lowest temperature the DCR, AP, and CT have been tested was 40 K, while the working temperature of the SiPMs equipped on LHe detectors is supposed to be near LHe temperature; so, we need to test these parameters near 4.5 K.
@@ III.3.1 DCR tests at 10 K
For the NUV-HD-Cryo type SiPMs under test, Fig. 11 shows the measured DCR is 0.01 Hz/mm2 for OV 10 V and 12 V at 10 K. Reference [27] measured the DCR for two types of FBK SiPMs: NUV-HD-SF and NUV-HD-LF. At RT, the typical DCR with 4 to 6 V OV is 100 kHz, while at 40 K, the DCR is 0.1 Hz/mm2 and 0.001 Hz/mm2 for the NUV-HD-SF and NUV-HD-LF SiPMs, respectively. Given that (a) the dimension of our SiPMs is 6 times that of the ones tested in the reference, and (b) the OV in our tests (10 to 12 V) is 6 volts higher than the paper (4 to 6 V), we consider our DCR results are consistent with the NUV-HD-LF ones measured in the paper at 40 K. We stick to a 10 V OV instead of a lower bias because it is the voltage the SiPMs can reach a required PDE criterion for LHe TPCs, 40%.
@@ III.3.2 AP tests at 10 K
The setup for the AP tests is almost the same as the DCR measurements (the LED was off either); the only difference is that the counter was replaced with a digitizer, the DAQBOX-4-125 model, from TOFTEK company in China. The digitizer has a 125 MSPS sampling rate, a 12-bit resolution, and an input range of -1 V to +1 V. It operates in an auto-trigger mode: whenever a signal’s amplitude crosses a programmable threshold, the timing crossing the threshold will be registered as the pulse’s time, T0. A user-definable time interval will be used to define a time window. Assuming the interval is set to be 1 second, the time window would be (T0, T0 + 1)s. And all of the pulses within the window will be registered. For each recorded pulse, the time is the difference between its threshold-crossing timing and the T0; the charge is the integration of the analog signal (by the digitizer). Fig. 12 shows the time and charge info of the signals recorded in 10 seconds. During data-taking, the threshold was set to be 40 mV 555The 40 mV was measured with a 5 amplifier connected upstream. So, it is equivalent to 8 mV in Fig. 6, which was directly measured on the preamplifier’s output., and the time window of each pulse was 2 s. The 2 s time window is wide enough for the 1.5 s width signals (Fig. 6) and narrow enough to exclude unnecessary noises. As shown in Fig. 12, there are 768 AP events and 4076 primary DCR and CT events. The AP ratio is calculated to be 768 / (768 + 4076) = 15.9%.
At 10 K, we also tested the AP under other bias voltages, as Fig. 13 shows. The measured AP rate at 10 K is (surprisingly) consistent with the results at 40 K [27] for the NUV-HD-LF type SiPMs in the sense that (a) the relationship between the bias voltage and the AP rate is the same: the higher the bias, the greater the AP; and (b) with the same bias voltage of OV 5 V or 6 V, the measured rate are both around 10%.
@@ III.3.3 CT tests at 10 K
As shown in Fig. 12, the CT and primary DCR events overlap on the time-charge plane. These two categories of events can not be distinguished in timing because the CT events intrinsically have the exact timing as the primary DCR ones. On the other hand, since the charge distribution does not feature individual peaks, as Fig. 7 shows, we cannot do the CT analysis as reference [27]. By assuming (a) the CT photons propagated as a branching Poisson process and (b) the signals produced in the SiPMs followed the Borel distribution, reference [52] modeled the CT events well for the SiPMs operated at RT. In this work however, we figured the CT out for the SiPMs near LHe temperature in an alternative way, as follows.
The data for the CT analysis was the same as for the charge analysis when the SiPMs was illuminated with weak light, as shown in Fig. 7. The CT analysis can be split into three steps: (a) making a charge fit between the “0 Phe” and “1 Phe” region only, as the red-on-green part of the curve in Fig. 7 shows; (b) based on the fit results in step (a), extrapolating the actual number of events of the histogram, N, under the (reasonable) assumption of no CT events included; (c) calculating the CT rate with the equation of (N - N) / N, where N is the charge histogram’s total events, which is 3000 in Fig. 7. We believe almost all of the (N - N) events are the CT events, while the AP ones are minimum because (i) although the oscilloscope screen has a window of 10 s and the SiPMs’ signal is relatively narrow, 1.5 s, as shown in Fig. 6, we have set a 2 s time window on the screen with cursors in data taking to make sure only the “true” signals (inside of the 2 s window) have been analyzed, while the “fake” events (outside of the window) have been excluded; and (ii) as shown in Fig. 12, all of the AP events have a 2 s delay from the primary DCR events. As a result, Fig. 7 would only have ignorable AP events, and (N - N) represents CT events.
Since the charge fit has already been introduced in detail in section III.2.1 above and step (c) is very straightforward, we mainly focus on step (b) here. We first select part of the histogram and the fit function, on which the equivalent relationship between the number of events of the partial histogram and the area of the partial fit function can be built. Then, we extrapolate the N with the relationship and the area of the total fit function. In the following, we intentionally ignore the units to simplify the analysis.
Only one of the 388 selected events was contributed from the “2 Phe” events (The event would be visible as long as the Y-axis is on a log scale). The area of the partial fit function is the integration between -1 and 0.3318, 42.52211.6260, which corresponds to the region being enclosed by the fit curve, the X-axis, and the X = 0.3318 line in Fig .14. The integral of 42.5221 and its uncertainty of 1.6260 were obtained from calling the fit function’s ROOT-inlined functions, “Integral ()” and “IntegralError ()”, respectively. Given the quality of the fit in Fig. 7, the area of 42.5221 and the 388 events can be considered equivalent. Since the area of the whole fit function (X from -1 to 8 of Fig. 7) is 314.8500 5.7451, the histogram should have 2872.9014 121.7230 “actual” events (without CT events interference), as Eq. (9) extrapolates,
As Fig. 7 indicates, the charge histogram has 3000 events in total, which means there are 127.0986 121.7230 ( = 3000 - (2872.9014 121.7230) ) events contributed from CT. The CT event ratio can be calculated as Eq. (10),
@@
which is similar to the measured CT for the NUV-HD-LF type of FBK SiPMs biased with OV = 3 V at 40 K, 6% [34]. Under higher bias voltages (OV = 4, 5, 6 V), the measured CT in the reference is between 10 and 16%. We measured the SiPMs’ CT being biassed with 5 to 11 V OV and calculated the uncertainties, as summarized in Fig. 15. In the plot, all of the CT rate values and their uncertainties were normalized to have 1.0 Phe. Further, a linear fit was applied to fit the CT rate data. P0 is the constant, and P1 is the slope. We saw the expected trend of the higher the bias, the greater the CT rate.

### v5

@@ III.2.1 PDE tests at 10 K
We implemented the photocurrent method to measure the PDE, as introduced in reference [Zappala16]. The PDE is calculated as Eq. (1),
where and are the measured current of the SiPMs with the light on and off, respectively; is the gain of the SiPMs; is the charge of an electron; stands for Excess Charge Factor, which represents the fake efficiency prompt contributed from AP and CT [Piemonte12, MPPCTechnicalNote21]; is the incident photons.
@@
where and represent the measured current for the bias voltage above or below the breakdown voltage in the dark, respectively; the definition of and are the same as Eq. (1); is the average number of photoelectrons per second in dark condition, which is equivalent to the average number of the Poisson statistical photoelectrons in a test [Otte06, Finocchiaro09, Eckert10, Yang14, Piemonte16, Zappala16]. As mentioned above in section III.1, the SiPMs’ DCR is as low as 1 Hz near LHe temperature in the dark, which leads to 0.1 pA current. While the measured current is 0.1 nA, essentially the detector system’s dark current. So, we illuminated the SiPMs with a weak pulsed light for the ECF tests to have a A scale current. Effectively, our testing condition of “weak light at LHe temperature” is equivalent to the conventional one of “dark at RT”.
@@
The SiPMs were illuminated with 405 nm or 530 nm light from an LED driven by a DG535 pulse generator. For PDE tests, the pulses from the DG535 are 4 V amplitude, 240 ns width, and 10 kHz frequency. Whenever the DG535 generated a pulse, it also triggered an 8 GHz bandwidth, 25 GSa/s sample rate oscilloscope to record the waveform showing on the screen. A typical single photoelectron (Phe) signal of the FBK SiPMs at 10 K is shown in Fig. 6. Given that (a) the effective time window of the pulses is 2.4e-3 s (= 240 ns 10 kHz) per second, and (b) the DCR is about 1 Hz near LHe temperature, the dark current events only have a chance of 0.24% (= 2.4e-3 s / 1 s) to show up in data. Therefore, the DCR will not smear the PDE results.
@@
As mentioned above, we must figure out the average number of photoelectrons to calculate PDE. We get it (mainly) through a fit on a charge histogram as shown in Fig. 7. The histogram was obtained by integrating thousands of analog signals in a 10 s window. Fig. 6 shows such a typical signal. During data-taking, the DG535 was tuned to generate weak enough light to produce 1.5 Phe in the SiPMs on average. The oscilloscope was triggered by the DG535. The signals were firstly recorded in the oscilloscope, then analyzed offline with ROOT [rootWebsite]-based scripts. A recorded raw signal often contained roughly 2000 data points, which were integrated directly (without baseline subtraction), then divided by a 50 ohm (the input resistance of the oscilloscope) to convert into charge. Next, we multiplied the charge with the system’s gain to get the number of Phe, as shown in Eq. (4),
@@
where the is the integral of raw signals, 50 is the input resistance of the oscilloscope, the is the gain of the SiPMs (often is the scale of 1.0 106) and the preamplifier combined, 1.6 10-19 is the charge of an electron. With the same setup and analysis, under the OV between 5 and 11 V, all of the charge histograms have shown explicit separated Phe peaks. Fig. 7 shows a typical charge distribution with OV = 6 V and a fit. The fit function is Gaussian convoluted Poisson, as shown in Eq. (7).
@@
The meaning of the parameters in the above equations are the following. “”, “”, and “” are the fit functions correspond to the pedestal (or 0 Phe), 1 to N Phe peaks, and the whole histogram, respectively. “” is the average number of the Poisson function. “” means the scale factor of the fit function on the Y-axis. “” indicates the shift on the X-axis, which is the analog signals’ baseline shift on the oscilloscope. “” corresponds to the scale factor on the X-axis. “” and “” are the 1- of the one Phe peak and the pedestal distribution, respectively.
@@
In Fig. 7, we limited our fit function on the less than 2 Phe part of the histogram to avoid smearing the fit by AP and CT events [Acerbi17, AcerbiGundacker19]. The average number of photoelectrons, 1.42 0.02 Phe, is the ND to be substituted in Eq. (3) for PDE calculation. As will be discussed further in III.3.3, the 1.42 0.02 is consistent with the ratio of the events in the “0 Phe” and “1 Phe” peaks assuming the events follow a Poisson distribution, therefore, justifying the reliability of the fit. The events in the “0 Phe” and “1 Phe” peaks are 672 and 963, respectively; and 672 / 963 Poisson (0, 1.42) / Poisson (1, 1.42), where Poisson (0, 1.42) and Poisson (1, 1.42) are the probabilities of 0 and 1 under a Poisson distribution with 1.42 on average, respectively. Other parameters we measured or calculated for the PDE test are shown in table 1. Submitting these parameters into Eq. (3) gives the PDE of (23.02 1.18)%. Following the same procedure, the PDE values measured with other bias voltages and 530 nm light can also be obtained, as shown in Fig. 8. From the plot, we can see that for a bias voltage of 10 V OV or more, the PDE is equal to or greater than 40%.
@@
| 1.420.02 | 4.770.09 | 0.120.07 | 2.640.02 | 0.020.01 | (6.830.31)E6 | 23.02 1.18 |
III.2.2 Gain tests at 10 K
We tested the SiPMs’ gain at 10 K with 405 nm light, as shown in Fig. 9. With the same bias voltage of OV = 5 V, the gain of the NUV-HD-Cryo SiPMs at 10 K was 0.23 106; while the gain of the NUV-HD-LF type of SiPMs at 40 K was 0.32 106 [Acerbi17, DarkSide20k17]. Our results are consistent with previous measurements, which observed the feature of gain decreasing with temperature. However, we identified a difference in the relationship between gain and overvoltage. Tests on the same type of SiPMs at liquid nitrogen temperature reported a linear relationship between 2 - 10 V OV [Acerbi17, Capasso20], a feature not observed in our tests (Fig. 9). Given that a non-linear relationship was also observed in an independent test on a different type of SiPMs at 100 mK [SiPMsTests100mk], we suspect the linear relationship may not hold for temperatures below 10 K, although a convincing model is still lacking. We will continue to investigate this issue in future tests.
@@ III.3 DCR, AP, and CT tests at 10 K
The dark current rate (DCR), after-pulse (AP), and cross-talk (CT) are critical parameters for SiPMs. Although these parameters were measured for two types of FBK SiPMs from RT down to 40 K with a different DAQ system, as shown in reference [Zappala16, Acerbi17], we must characterize them near LHe temperature again for the following reasons. (a) The SiPMs under our tests are not the same model as the references according to our discussion with FBK engineers, though it is similar to the NUV-HD-LF one. (b) The dimension of our SiPMs is 11.7 7.9 mm2, which is a factor of six greater than the ones tested in reference [Zappala16, Acerbi17], 4.0 4.0 mm2; the cell pitch of our SiPMs is 30 30 m2, which is also bigger than the 25 25 m2 one in the references. (c) For FBK SiPMs, the lowest temperature the DCR, AP, and CT have been tested was 40 K, while the working temperature of the SiPMs equipped on LHe detectors is supposed to be near LHe temperature; so, we need to test these parameters near 4.5 K.
@@ III.3.1 DCR tests at 10 K
For the NUV-HD-Cryo type SiPMs under test, Fig. 10 shows the measured DCR is 0.01 Hz/mm2 for OV 10 V and 12 V at 10 K. Reference [Acerbi17] measured the DCR for two types of FBK SiPMs: NUV-HD-SF and NUV-HD-LF. At RT, the typical DCR with 4 to 6 V OV is 100 kHz, while at 40 K, the DCR is 0.1 Hz/mm2 and 0.001 Hz/mm2 for the NUV-HD-SF and NUV-HD-LF SiPMs, respectively. Given that (a) the dimension of our SiPMs is 6 times that of the ones tested in the reference, and (b) the OV in our tests (10 to 12 V) is 6 volts higher than the paper (4 to 6 V), we consider our DCR results are consistent with the NUV-HD-LF ones measured in the paper at 40 K. We stick to a 10 V OV instead of a lower bias because it is the voltage the SiPMs can reach a required PDE criterion for LHe TPCs, 40%.
@@ III.3.2 AP tests at 10 K
The setup for the AP tests is almost the same as the DCR measurements (the LED was off either); the only difference is that the counter was replaced with a digitizer, the DAQBOX-4-125 model, from TOFTEK company in China. The digitizer has a 125 MSPS sampling rate, a 12-bit resolution, and an input range of -1 V to +1 V. It operates in an auto-trigger mode: whenever a signal’s amplitude crosses a programmable threshold, the timing crossing the threshold will be registered as the pulse’s time, T0. A user-definable time interval will be used to define a time window. Assuming the interval is set to be 1 second, the time window would be (T0, T0 + 1)s. And all of the pulses within the window will be registered. For each recorded pulse, the time is the difference between its threshold-crossing timing and the T0; the charge is the integration of the analog signal (by the digitizer). Fig. 11 shows the time and charge info of the signals recorded in 10 seconds. During data-taking, the threshold was set to be 40 mV 555The 40 mV was measured with a 5 amplifier connected upstream. So, it is equivalent to 8 mV on the preamplifier’s output., and the time window of each pulse was 2 s. The 2 s time window is wide enough for the 1.5 s width signals 666The preamplifier we implemented in the AP tests is different from the one being utilized for PDE and CT tests. Among other difference, the former has a 1.5 s integration time, while the latter is s, as shown in Fig. 6. and narrow enough to exclude unnecessary noises. As shown in Fig. 11, there are 768 AP events and 4076 primary DCR and CT events. The AP ratio is calculated to be 768 / (768 + 4076) = 15.9%.
At 10 K, we also tested the AP under other bias voltages, as Fig. 12 shows. The measured AP rate at 10 K is (surprisingly) consistent with the results at 40 K [Acerbi17] for the NUV-HD-LF type SiPMs in the sense that (a) the relationship between the bias voltage and the AP rate is the same: the higher the bias, the greater the AP; and (b) with the same bias voltage of OV 5 V or 6 V, the measured rate are both around 10%.
@@ III.3.3 CT tests at 10 K
The data for the CT analysis was the same as for the charge analysis when the SiPMs was illuminated with weak light, as shown in Fig. 7. The CT analysis can be split into three steps: (a) Figuring out the average number of photoelectrons, , based on the number of events in “0 Phe” and “1 Phe”, which correspond to the first and second peaks in Fig. 7, respectively; (b) Extrapolating the expected number of events in the histogram, N, under the reasonable assumption of no CT events included; (c) calculating the CT rate with the equation of (N - N) / N, where N is the charge histogram’s total events, which is 3000 in Fig. 7. In the figure, the events in the ‘0 Phe” and “1 Phe” peaks are 672 and 963, respectively. Assuming these events follow a Poisson distribution, the of the histogram would be 1.43 since 672 / 963 Poisson (0, 1.43) / Poisson (1, 1.43), while the fit result is 1.42 0.02, perfectly lines up with the calculation based on the events in the two peaks. In addition, since all the charge histograms we measured with OV between 5 and 11 V have explicitly separated Phe peaks, which means there are no smearing events across the peaks. As a result, for the CT analysis, we rely on the events in the peaks (instead of the fit results).
According to the events in the “0 Phe” (672) and “1 Phe” (963) peaks, the average number Phe in Fig. 7 is 1.43 since 672 / 963 Poisson (0, 1.43) / Poisson (1, 1.43), so the expected events, N, can be obtained as the following Eq.(8) provided no smearing events in the 2+ peaks.
The observed events in all the 2 + peaks in Fig. 7 are 1362. So the discrepancy, 1362 - 1177 = 185, is the number of CT events. And the preliminary CT ratio is 185 / 3000 = 0.0615 = 6.15%. Since (a) a CT event happens only when one photon emits, and (b) the mean Phe are not exactly the same in each dataset with different OV values because the LED’s light intensity decreases along with the running time around 10 K according to our tests, so we normalized the CT ratio corresponding to 1 Phe by dividing the preliminary one with to get the final value, 5.15%. In uncertainty estimation, we simplified it by only considering the errors in the final step of calculation, i.e, the preliminary ratio. The errors for 185 and 3000 are (= 13.59) and (= 54.77), respectively. We then propagate the errors according to the calculation ( (185 / 3000) / ) to get the final uncertainty, 6.33%. And the final CT ratio is 5.15% 6.33%.
@@
We measured the SiPMs’ CT being biassed with 5 to 11 V OV and calculated the uncertainties, as summarized in Fig. 13 777Be aware please, the CT analysis we did here is different from others, eg, Ref. [Eckert10].. In the plot, all of the CT rate values and their uncertainties were normalized to have 1.0 Phe. Further, a linear fit was applied to fit the CT rate data. P0 is the constant, and P1 is the slope. We saw the expected trend of the higher the bias, the greater the CT rate.